# Modern Neural Networks for Small Tabular Datasets: The New Default for Field-Scale Digital Soil Mapping?

### v1

@@ Abstract
In the field of pedometrics, tabular machine learning is the predominant method for predicting soil properties from remote and proximal soil sensing data, forming a central component of digital soil mapping. At the field-scale, this predictive soil modeling (PSM) task is typically constrained by small training sample sizes and high feature-to-sample ratios in soil spectroscopy. Traditionally, these conditions have proven challenging for conventional deep learning methods. Classical machine learning algorithms, particularly tree-based models like Random Forest and linear models such as Partial Least Squares Regression, have long been the default choice for field-scale PSM. Recent advances in artificial neural networks (ANN) for tabular data challenge this view, yet their suitability for field-scale PSM has not been proven. We introduce a comprehensive benchmark that evaluates state-of-the-art ANN architectures, including the latest multilayer perceptron (MLP)-based models (TabM, RealMLP), attention-based transformer variants (FT-Transformer, ExcelFormer, T2G-Former, AMFormer), retrieval-augmented approaches (TabR, ModernNCA), and an in-context learning foundation model (TabPFN). Our evaluation encompasses 31 field- and farm-scale datasets containing 30 to 460 samples and three critical soil properties: soil organic matter or soil organic carbon, pH, and clay content. Our results reveal that modern ANNs consistently outperform classical methods on the majority of tasks, demonstrating that deep learning has matured sufficiently to overcome the long-standing dominance of classical machine learning for PSM. Notably, TabPFN delivers the strongest overall performance, showing robustness across varying conditions. We therefore recommend the adoption of modern ANNs for field-scale PSM and propose TabPFN as the new default choice in the toolkit of every pedometrician.
@@ 1 Introduction
Soil maps are relevant for a range of environmental and agricultural applications (Keesstra et al., 2016). Especially at the field-scale, they play a crucial role in supporting high crop yields while considering negative environmental impacts (Gebbers and Adamchuk, 2010; Viscarra Rossel and Bouma, 2016). Sensors enable rapid and cost-effective generation of high-resolution soil data needed for creating these soil maps. This includes off-site sensing in the laboratory or direct in-situ measurements in field-conditions through proximal-soil sensing (PSS) or remote sensing (RS) (Viscarra Rossel and Bouma, 2016). However, these sensors do not directly measure actual soil properties relevant for agriculture but proxies related to them (Gebbers, 2019). The translation of sensor signals into useful soil property estimates requires a set of soil samples to train a site-specific prediction model. This predictive soil modeling (PSM) framework falls in the field of pedometrics and is used for digital soil mapping (McBratney et al., 2003, 2019).
@@
Tabular machine learning (ML) has emerged as the predominant approach in contemporary PSM  (Heuvelink and Webster, 2022). However, the application of ML for field-scale PSM faces specific challenges, distinguishing it from other ML domains. These challenges include high noise, complexity of soil as a medium, variable measurement conditions, differences in measurement footprints and spatial autocorrelation. Most importantly, training sample sizes are generally rather low as soil sampling and soil analyses in laboratory rapidly becomes cost- and labor-prohibitive  (Söderström et al., 2016). This creates a fundamental tension: while sample sets cannot be too small as this would result in poorly fitted prediction models, economic considerations constrain the size of training sample sets (Schmidinger et al., 2024). Additionally, the feature dimensionality in PSM can vary drastically from very low dimensional data generated by common in-situ PSSs like electrical conductivity sensors to very high dimensional data obtained from mid- and near-infrared spectroscopy. In soil spectroscopy, the feature dimensionality usually exceeds the number of samples, making modeling prone to overfitting (dos Santos et al., 2023; Wang and Wang, 2022; Wang et al., 2022), a challenge known as the curse of dimensionality (Bellman et al., 1957).
@@
These dataset characteristics favored classical ML approaches like tree-based algorithms and linear models, which became the default choices for PSM (Wadoux et al., 2020; Barra et al., 2021; Ding et al., 2025). In particular, Random Forest dominates due to its straightforward application, good performance with default hyperparameters (Probst et al., 2019), fast training times, and general effectiveness even with low training sample sizes (Ma et al., 2020; Schmidinger et al., 2024). Recent reviews consistently report Random Forest as the most frequently used algorithm for PSM (Ding et al., 2025; Wadoux et al., 2020). In soil spectroscopy applications specifically, linear models, usually combined with linear feature transformations such as principal component analysis (PCA) or partial least squares, have established themselves as fundamental methods (Barra et al., 2021). They demonstrate robust performances in contemporary spectroscopy studies, even when evaluated alongside popular ML algorithms such as Random Forest or Artificial Neural Networks (ANN) (e.g. Xue et al., 2023; Schmidinger et al., 2025).
@@
Deep learning applications in the form of various ANN architectures, notably Multilayer Perceptrons (MLP) and Convolutional Neural Networks, have attracted increasing interest for PSM (Wadoux, 2025), but have mostly shown success for contexts where abundant training data is available. This includes large-area soil mapping at a regional up to continental level utilizing RS (e.g. Yang et al., 2021) or vast spectral libraries (e.g. Ng et al., 2019). However, this success did not extend to field- or smaller regional-scale applications where sample sizes are commonly lower, with ANNs showing relatively poor performances (e.g. Xue et al., 2023; Oukhattar et al., 2025). Moreover, even in large-area applications, deep learning did not always show improvements over classical methods (e.g. Sun et al., 2023). These limitations, particularly in smaller area high resolution soil mapping, are further emphasized by the methodological comparison of Khaledian and Miller (2020), who argue that classical ML methods provide more stable results than ANNs in cases of limited training data. In fact, Khaledian and Miller (2020) explicitly recommend using classical ML algorithms when working with fewer than 1,000 soil samples. Overall, the effectiveness of deep learning for PSM remains uncertain, and it has not replaced classical ML methods.
@@
The historical dominance of classical ML for PSM aligns with broader trends in the tabular data domain, where classical ML approaches have maintained a competitive edge over ANNs until recent years. Following deep learning’s widespread success in computer vision and natural language processing, early attempts to apply ANNs to tabular data generated considerable optimism, with several studies reporting promising results (Arik and Pfister, 2021; Katzir et al., 2020). However, subsequent benchmarking studies revealed that classical ML, particularly tree-based ensemble methods like gradient-boosted decision trees (GBDTs), consistently outperformed these early deep learning approaches (Shwartz-Ziv and Armon, 2022). In fact, Shwartz-Ziv and Armon (2022) observed that the early claims of ANN superiority were largely driven by limited benchmarking practices, with studies relying on favorable results from a narrow selection of datasets, failing to generalize across broader evaluations. This was subsequently confirmed by other comprehensive tabular benchmarks (Grinsztajn et al., 2022; McElfresh et al., 2023). Notably, these benchmarks confirmed classical ML superiority in conditions with limited sample sizes and high dimensionality. McElfresh et al. (2023) observed that the performance gap between tree-based models and ANNs widened considerably as the dataset size decreased, specifically noting that ANNs performed comparatively worse when the feature-to-sample ratio became larger.
@@
Recent developments in deep learning for tabular data challenge this perspective. Over the past year, multiple ANN approaches claiming to surpass classical ML methods have been proposed for tabular data (Hollmann et al., 2025; Holzmüller et al., 2024; Gorishniy et al., 2025; Ye et al., 2025c; Cheng et al., 2024). In contrast to previous deep learning models, recent tabular benchmarks also support this shift (Erickson et al., 2025; Ye et al., 2025b). These advances extend far beyond incremental improvements to the classically used ANN architectures like MLPs. Entirely new tabular ANN approaches have emerged, including attention-based models, retrieval-based approaches, and in-context learning foundation models (Ye et al., 2025a). We refer to these recent architectural innovations, from the past year, collectively as modern ANN approaches throughout this manuscript. These modern ANNs hold particular promise for addressing PSM-specific challenges. In-context learning models such as TabPFN (Hollmann et al., 2025) claim to achieve particularly strong performance on small datasets, making them potentially strong choices for the small-area PSM domain. Retrieval architectures such as ModernNCA (Ye et al., 2025c) exploit sample neighborhood relationships, and may offer unique advantages for soil property prediction given the spatially correlated nature of soil data.
@@
Preliminary ML benchmarking studies showed impressive performance on public domain-independent benchmarks with these modern ANN approaches  (Ye et al., 2025a; Erickson et al., 2025; Hollmann et al., 2025; Ye et al., 2025b). However, these studies did not adequately represent the specific constraints of field-scale PSM, as their dataset properties differed substantially from those encountered in field-scale PSM (see Appendix C). For example, datasets in the benchmark of Erickson et al. (2025) and Ye et al. (2025a) only included datasets with at least 500 samples. While this sample size is considered small in the broader ML context, it still largely exceeds the maximum sample size commonly available in field-scale PSM (Schmidinger et al., 2025). Additionally, the aspect of high-dimensionality and unfavorable feature-to-sample ratios is often dismissed or marginally discussed. Consequently, there remains a crucial need to examine the applicability and efficacy of these advanced deep learning paradigms in the context of the inherent constraints of field-scale PSM.
@@
Applying and comparing these promising architectures in field-scale PSM is non-trivial. As the broader tabular data literature already warns (Shwartz-Ziv and Armon, 2022), proper evaluation of the new deep learning methods is itself a significant methodological challenge, and inadequate benchmarking practices can substantially affect conclusions about their relative performance. Methodological problems of previous tabular benchmarking studies have been largely overlooked in pedometrics, leading to critical limitations that may produce incomplete or misleading results (Schmidinger et al., 2025). Recent work by Schmidinger et al. (2025) revealed that over 95% of PSM benchmarking studies relied on a single dataset, with the largest number of datasets used in any reviewed study being only three. As demonstrated by Shwartz-Ziv and Armon (2022), reliance on a single dataset may inadvertently result in overinterpretation of incidental findings, obscuring the actual strengths and limitations of deep learning methods for PSM. Other confounding factors, like the choice of hyperparameters, can also introduce significant bias (Nießl et al., 2022). In some PSM studies, hyperparameters of ANNs were carefully selected, whereas the competing tree-based models have not been tuned or no information has been given (e.g. Ng et al., 2019; Yang et al., 2021). Given that fewer than 10% of PSM benchmarking studies provided open datasets and less than 5% shared their code (Schmidinger et al., 2025), it is usually not possible to reproduce and verify the results. These shortcomings highlight the critical need for comprehensive benchmarking studies in PSM that systematically assess novel methods across multiple diverse datasets in a fair and rigorous way. Different training regimes, parameter optimization approaches, validation strategies, and other methodological choices must be implemented consistently to ensure fair comparisons (Nießl et al., 2022).
@@
In this work, we benchmark modern deep learning methods that have shown promise in tabular data applications but remain unexplored in the domain of PSM. Our study makes several key contributions. We are the first to systematically introduce and assess state-of-the-art deep learning methods for PSM at the field-scale with its distinct dataset challenges regarding sample size and feature dimensionality. Second, we establish a fair evaluation framework with consistent hyperparameter optimization, training regimes, and validation strategies across all methods. Third, we conduct experiments across multiple diverse field-scale soil datasets, avoiding the single-dataset evaluation dominating in previous benchmarking studies in pedometrics. Finally, we ensure complete reproducibility by making all datasets, code, and experimental configurations publicly available (see Code and data availability). Through this analysis, we aim to provide empirical evidence on whether these advanced deep learning methods can overcome the longstanding dominance of classical ML in pedometric applications, particularly under the small-sample constraints typical for farm-scale PSM.
@@ 2.1 Datasets
Figure 1: Global distribution of the 31 field- and farm-scale datasets used in this study. Background map is based on Homolosine projection. The datasets span multiple continents and diverse agricultural contexts, including locations across North and South America, Europe, Asia, and Australia.
@@
We utilized Precision Liming Soil Datasets (LimeSoDa), a collection of 31 field- to farm-scale datasets spanning multiple countries and diverse agricultural contexts, representing a broad spectrum of soil mapping scenarios encountered in precision agriculture (Schmidinger et al., 2025). Their global distribution is illustrated in Figure 1. Each dataset contains three target soil properties: soil organic matter (SOM) or soil organic carbon (SOC), pH, and clay content, yielding a total of 93 regression tasks for model evaluation. Individual dataset sizes range from 30 to 460 samples, providing varied data scenarios typical of real-world PSM projects for small-area soil mapping. All datasets are openly accessible and have been standardized in tabular format to facilitate reproducible research through the LimeSoDa repository (see Code and data availability).
@@
Features are dataset-specific and were obtained by different sensing technologies, including laboratory-based spectroscopy, in-situ PSS, and RS. This includes a variety of common sensor modalities, such as apparent electrical resistivity, gamma-ray spectrometry, ion selective electrodes, digital elevation models, multispectral RS data, vegetation indices, and X-ray fluorescence derived elemental concentrations. Several datasets incorporate high-dimensional optical spectroscopy data from visible and near-infrared (vis-NIR), near-infrared (NIR), and mid-infrared (MIR) spectrometers. These measurements result in feature sets with up to 2,489 variables, as the reflectance is typically recorded across numerous continuous wavelength bands. There are no categorical features present across any of the datasets. See Appendix C for comparisons to feature sizes and feature-to-sample ratios reported in prior benchmarks.
Given the computational challenges posed by high-dimensional spectral data, we categorized the datasets into two groups based on feature characteristics and dimensionality for our analysis (Figure 2).
@@
The first group, referred to as "High-Dimensional", includes datasets with feature-to-sample ratios >>1, containing vis-NIR, NIR, or MIR spectroscopy features. These datasets require dimensionality reduction because the unfavorable feature-to-sample ratio would otherwise lead to severe overfitting (Schmidinger et al., 2025). Dimensionality reduction was performed during preprocessing as described in Section 2.4.3.
The second group, referred to as "Low-Dimensional", includes datasets with lower-dimensional PSS features (e.g. apparent electrical resistivity) or RS features (e.g. vegetation indices). All of these datasets have a feature-to-sample ratio <<1.
@@
Figure 2: Distribution of datasets by number of features and samples (log scale), highlighting feature-to-sample ratios. Each point represents one dataset, with two groups shown: Low-Dimensional datasets (lower feature counts) and High-Dimensional datasets (higher feature counts due to spectral features).

### v2

@@ Abstract
In the field of pedometrics, tabular machine learning is the predominant method for soil property prediction from remote and proximal soil sensing data, forming a central component of Digital Soil Mapping (DSM). At the field-scale, this pedometric modeling task is typically constrained by small training sample sizes and high feature-to-sample ratios, particularly when using spectroscopic data. Traditionally, these conditions have proven challenging for conventional deep learning methods. Classical machine learning algorithms, particularly tree-based models like Random Forest and linear models such as Partial Least Squares Regression, have long been the default choice for pedometric modeling within DSM. Recent advances in artificial neural networks (ANN) for tabular data challenge this view, yet their suitability for field-scale DSM has not been proven. We introduce a comprehensive benchmark that evaluates state-of-the-art ANN architectures, including the latest multilayer perceptron (MLP)-based models (TabM, RealMLP), attention-based transformer variants (FT-Transformer, ExcelFormer, T2G-Former, AMFormer), retrieval-augmented approaches (TabR, ModernNCA), and an in-context learning foundation model (TabPFN). Our evaluation encompasses 31 field- and farm-scale datasets containing 30–460 soil samples and three critical soil properties: soil organic matter or soil organic carbon, pH, and clay content. Our results reveal that modern ANNs consistently outperform classical methods on the majority of tasks, demonstrating that deep learning has matured sufficiently to overcome the long-standing dominance of classical machine learning for soil property prediction. Notably, TabPFN delivers the strongest overall performance, showing robustness across varying conditions. We therefore recommend the adoption of modern ANNs for field-scale DSM and propose TabPFN as the new default choice in the toolkit of every pedometrician.
@@
[…]
Notice: The published version of this article is available in European Journal of Soil Science at: https://doi.org/10.1111/ejss.70299
@@ 1 Introduction
Soil maps are relevant for a range of environmental and agricultural applications (Keesstra et al., 2016). Especially at the field-scale, they play a crucial role in supporting high crop yields while considering negative environmental impacts (Gebbers and Adamchuk, 2010; Viscarra Rossel and Bouma, 2016). Sensors enable rapid and cost-effective generation of high-resolution soil data needed for creating these soil maps. This includes off-site sensing in the laboratory or direct in-situ measurements in field-conditions through proximal-soil sensing (PSS) or remote sensing (RS) (Viscarra Rossel and Bouma, 2016). However, these sensors do not directly measure actual soil properties relevant for agriculture but proxies related to them (Gebbers, 2019). The translation of sensor signals into useful soil property estimates requires a set of soil samples to train a site-specific prediction model. This pedometric modeling step is central to Digital Soil Mapping (DSM), the creation of spatially referenced soil information through quantitative models that infer the spatial variations of soil properties from soil observations and high-resolution environmental covariates (McBratney et al., 2003, 2019).
@@
Tabular machine learning (ML) has emerged as the predominant approach for pedometric modeling in contemporary DSM  (Heuvelink and Webster, 2022). However, the application of ML for field-scale DSM faces specific challenges, distinguishing it from other ML domains. These challenges include high noise, complexity of soil as a medium, variable measurement conditions, differences in measurement footprints and spatial autocorrelation. Most importantly, training sample sizes are generally rather low as soil sampling and soil analyses in laboratory rapidly becomes cost- and labor-prohibitive  (Söderström et al., 2016). This creates a fundamental tension: while datasets cannot be too small as this would result in poorly fitted prediction models, economic considerations constrain the size of training sample sets (Schmidinger et al., 2024). Additionally, the feature dimensionality in DSM can vary drastically from very low dimensional data generated by common in-situ PSSs like electrical conductivity sensors to very high dimensional data obtained from mid- and near-infrared spectroscopy. In soil spectroscopy, the feature dimensionality usually exceeds the number of soil samples, making modeling prone to overfitting (dos Santos et al., 2023; Wang and Wang, 2022; Wang et al., 2022), a challenge known as the curse of dimensionality (Bellman et al., 1957).
@@
These dataset characteristics favored classical ML approaches like tree-based algorithms and linear models, which became the default choices for pedometric modeling within DSM (Wadoux et al., 2020a; Barra et al., 2021; Ding et al., 2025). In particular, Random Forest dominates due to its straightforward application, good performance with default hyperparameters (Probst et al., 2019), fast training times, and general effectiveness even with low training sample sizes (Ma et al., 2020; Schmidinger et al., 2024). Recent reviews consistently report Random Forest as the most frequently used algorithm for soil property prediction in DSM (Ding et al., 2025; Wadoux et al., 2020a). In soil spectroscopy applications specifically, linear models, usually combined with linear feature transformations such as principal component analysis (PCA) or partial least squares, have established themselves as fundamental methods (Barra et al., 2021). They demonstrate robust performance in contemporary spectroscopy studies, even when evaluated alongside popular ML algorithms such as Random Forest or Artificial Neural Networks (ANN) (e.g., Xue et al., 2023; Schmidinger et al., 2025).
@@
Deep learning applications in the form of various ANN architectures, notably Multilayer Perceptrons (MLP) and Convolutional Neural Networks, have attracted increasing interest for DSM (Wadoux, 2025), but have mostly shown success for contexts where abundant training data is available. This includes large-area soil mapping at a regional up to continental level utilizing RS (e.g., Yang et al., 2021) or vast spectral libraries (e.g., Ng et al., 2019). However, this success did not extend to field- or smaller regional-scale applications where sample sizes are commonly lower, with ANNs showing relatively poor performances (e.g., Xue et al., 2023; Oukhattar et al., 2025). Moreover, even in large-area applications, deep learning did not always show improvements over classical methods (e.g., Sun et al., 2023). These limitations, particularly in smaller area high resolution soil mapping, are further emphasized by the methodological comparison of Khaledian and Miller (2020), who argue that classical ML methods provide more stable results than ANNs in cases of limited training data. In fact, Khaledian and Miller (2020) explicitly recommend using classical ML algorithms when working with fewer than 1,000 soil samples. Overall, due to the uncertainty surrounding the effectiveness of deep learning for pedometric modeling in DSM, ANNs have not replaced classical ML approaches.
@@
The historical dominance of classical ML in DSM aligns with broader trends in the tabular data domain, where classical ML approaches have maintained a competitive edge over ANNs until recent years. Following deep learning’s widespread success in computer vision and natural language processing, early attempts to apply ANNs to tabular data generated considerable optimism, with several studies reporting promising results (Arik and Pfister, 2021; Katzir et al., 2020). However, subsequent benchmarking studies revealed that classical ML, particularly tree-based ensemble methods like gradient-boosted decision trees (GBDTs), consistently outperformed these early deep learning approaches (Shwartz-Ziv and Armon, 2022). In fact, Shwartz-Ziv and Armon (2022) observed that the early claims of ANN superiority were largely driven by limited benchmarking practices, with studies relying on favorable results from a narrow selection of datasets, failing to generalize across broader evaluations. This was subsequently confirmed by other comprehensive tabular benchmarks (Grinsztajn et al., 2022; McElfresh et al., 2023). Notably, these benchmarks confirmed classical ML superiority in conditions with limited sample sizes and high dimensionality. McElfresh et al. (2023) observed that the performance gap between tree-based models and ANNs widened considerably as the dataset size decreased, specifically noting that ANNs performed comparatively worse when the feature-to-sample ratio became larger.
@@
Recent developments in deep learning for tabular data challenge this perspective. Over the past two years, multiple ANN approaches claiming to surpass classical ML methods have been proposed for tabular data (Hollmann et al., 2025; Holzmüller et al., 2024; Gorishniy et al., 2025; Ye et al., 2025c; Cheng et al., 2024). In contrast to previous deep learning models, recent tabular benchmarks also support this shift (Erickson et al., 2025; Ye et al., 2025b). These advances extend far beyond incremental improvements to the classically used ANN architectures like MLPs. Entirely new tabular ANN approaches have emerged, including attention-based models, retrieval-based approaches, and in-context learning foundation models (Ye et al., 2025a). We refer to these recent architectural innovations, from the past year, collectively as modern ANN approaches throughout this manuscript. These modern ANNs hold particular promise for addressing pedometric modeling challenges in DSM. In-context learning models such as TabPFN, a Tabular Prior-data Fitted Network (Hollmann et al., 2025), are trained once on synthetic data to make predictions on new datasets from a context set of labeled examples provided in the input. They claim to achieve particularly strong performance on small datasets, making them potentially strong choices for the small-area DSM applications. Retrieval architectures such as ModernNCA (Ye et al., 2025c) exploit sample neighborhood relationships, and may offer unique advantages for soil property prediction given the spatially correlated nature of soil data.
@@
Preliminary ML benchmarking studies showed impressive performance on public domain-independent benchmarks with these modern ANN approaches  (Ye et al., 2025a; Erickson et al., 2025; Hollmann et al., 2025; Ye et al., 2025b). However, these studies did not adequately represent the specific constraints of field-scale DSM, as their dataset properties differed substantially from those encountered in field-scale DSM (see Table S1 in the Supporting Information E). For example, datasets in the benchmark of Erickson et al. (2025) and Ye et al. (2025a) only included datasets with at least 500 observations. While this sample size is considered small in the broader ML context, it still largely exceeds the maximum sample size commonly available in field-scale DSM (Schmidinger et al., 2025). Additionally, the aspect of high-dimensionality and unfavorable feature-to-sample ratios is often dismissed or marginally discussed. Consequently, there remains a crucial need to examine the applicability and efficacy of these advanced deep learning paradigms in the context of the inherent constraints of field-scale DSM.
@@
Applying and comparing these promising architectures in field-scale DSM is non-trivial. As the broader tabular data literature already warns (Shwartz-Ziv and Armon, 2022), proper evaluation of the new deep learning methods is itself a significant methodological challenge, and inadequate benchmarking practices can substantially affect conclusions about their relative performance. Methodological problems of previous tabular benchmarking studies have been largely overlooked in pedometrics, leading to critical limitations that may produce incomplete or misleading results (Schmidinger et al., 2025). Recent work by Schmidinger et al. (2025) revealed that over 95% of DSM benchmarking studies relied on a single dataset, with the largest number of datasets used in any reviewed study being only three. As demonstrated by Shwartz-Ziv and Armon (2022), reliance on a single dataset may inadvertently result in overinterpretation of incidental findings, obscuring the actual strengths and limitations of deep learning methods for pedometric modeling within DSM. Other confounding factors, like the choice of hyperparameters, can also introduce significant bias (Nießl et al., 2022). In some DSM studies, hyperparameters of ANNs were carefully selected, whereas the competing tree-based models have not been tuned or no information regarding their optimization was provided (e.g., Ng et al., 2019; Yang et al., 2021). Given that fewer than 10% of DSM benchmarking studies provided open datasets and less than 5% shared their code (Schmidinger et al., 2025), it is usually not possible to reproduce and verify the results. These shortcomings highlight the critical need for comprehensive benchmarking studies in DSM that systematically assess novel soil property prediction methods across multiple diverse datasets in a fair and rigorous way. Different training regimes, parameter optimization approaches, validation strategies, and other methodological choices must be implemented consistently to ensure fair comparisons (Nießl et al., 2022).
@@
In this work, we benchmark modern deep learning methods that have shown promise in tabular data applications but remain unexplored in the domain of pedometrics and DSM. Our study makes several key contributions. We are the first to systematically introduce and assess state-of-the-art deep learning methods for DSM at the field-scale with its distinct dataset challenges regarding sample size and feature dimensionality. Second, we establish a fair evaluation framework with consistent hyperparameter optimization, training regimes, and validation strategies across all methods. Third, we conduct experiments across multiple diverse field-scale soil datasets, avoiding the single-dataset evaluation dominating in previous benchmarking studies in pedometrics. Finally, we ensure complete reproducibility by making all datasets, code, and experimental configurations publicly available (see Code and data availability). Through this analysis, we aim to provide empirical evidence on whether these advanced deep learning methods can overcome the longstanding dominance of classical ML in pedometric applications, particularly under the small-sample constraints typical for farm-scale DSM.
@@ 2.1 Datasets
Figure 1: Global distribution of the 31 field- and farm-scale datasets used in this study. Background map is based on Homolosine projection. The datasets span multiple continents and diverse agricultural contexts, including locations across North and South America, Europe, Asia, and Australia. Base map source: Natural Earth.
@@
We utilized Precision Liming Soil Datasets (LimeSoDa), a collection of 31 field- to farm-scale datasets spanning multiple countries and diverse agricultural contexts, representing a broad spectrum of soil mapping scenarios encountered in precision agriculture (Schmidinger et al., 2025). Their global distribution is illustrated in Figure 1. Each dataset contains three target soil properties: soil organic matter (SOM) or soil organic carbon (SOC), pH, and clay content, yielding a total of 93 regression tasks for model evaluation. Individual dataset sizes range from 30 to 460 soil samples, providing varied data scenarios typical of real-world DSM projects for small-area soil mapping. All datasets are openly accessible and have been standardized in tabular format to facilitate reproducible research through the LimeSoDa repository (see Code and data availability).
@@
Features are dataset-specific and were obtained by different sensing technologies, including laboratory-based spectroscopy, in-situ PSS, and RS. This includes a variety of common sensor modalities, such as apparent electrical resistivity, gamma-ray spectrometry, ion selective electrodes, digital elevation models, multispectral RS data, vegetation indices, and X-ray fluorescence derived elemental concentrations. Several datasets incorporate high-dimensional optical spectroscopy data from visible and near-infrared (vis-NIR), near-infrared (NIR), and mid-infrared (MIR) spectrometers. These measurements result in feature sets with up to 2,489 variables, as the reflectance is typically recorded across numerous continuous wavelength bands. There are no categorical features present across any of the datasets. See Table S1 in the Supporting Information E for comparisons to feature sizes and feature-to-sample ratios reported in prior benchmarks.
@@
Given the computational challenges posed by high-dimensional spectral data, we categorized the datasets into two groups based on feature characteristics and dimensionality for our analysis (Figure 2). The first group, referred to as "High-Dimensional", includes datasets with feature-to-sample ratios >>1, containing vis-NIR, NIR, or MIR spectroscopy features. These datasets require dimensionality reduction because the unfavorable feature-to-sample ratio would otherwise lead to severe overfitting (Schmidinger et al., 2025). Dimensionality reduction was performed during preprocessing as described in Section 2.4.3. The second group, referred to as "Low-Dimensional", includes datasets with lower-dimensional PSS features (e.g., apparent electrical resistivity) or RS features (e.g., vegetation indices). All of these datasets have a feature-to-sample ratio <<1.
@@
Figure 2: Distribution of datasets by number of features and soil samples (log scale), highlighting feature-to-sample ratios. Each point represents one dataset, with two groups shown: Low-Dimensional datasets (lower feature counts) and High-Dimensional datasets (higher feature counts due to spectral features).